2 If f(x,y) = tan-1 xy then the value f(0.9, -1.2) is
Answer #1

f (x, y) = tant (ay) Now, f (0.9,-1.2) = tant (0.9+(-1.2)) = tant (-1.08) in degree 2-47-20259816 (approx.) =-82384075 (radiIf you have any doubt in this method or understanding, please comment and I will try my best to explore this to you. Thank you.